Mark watched an ant crawl through an arc of 40° along the rim of his melon, which was cut in half. If the radius of the melon was 7 inches, how far did the the ant crawl
Answer by rothauserc(4718)

a circle has 360 degrees and 40 degrees is 40/360 or 1/9 of a circle
circumference of a circle is pi*d where d is the denominator
d = 7*2 = 14, therefore
distance ant crawls = pi*14*(1/9) = 4.886921906 inches